# spacing=nonfrench

### v1

@@ Abstract
The limited generalization of coreference resolution (CR) models has been a major bottleneck in the task’s broad application. Prior work has identified annotation differences, especially for mention detection, as one of the main reasons for the generalization gap and proposed using additional annotated target domain data. Rather than relying on this additional annotation, we propose an alternative formulation of the CR task, Major Entity Identification (MEI), where we: (a) assume the target entities to be specified in the input, and (b) limit the task to only the frequent entities. Through extensive experiments, we demonstrate that MEI models generalize well across domains on multiple datasets with supervised models and LLM-based few-shot prompting. Additionally, the MEI task fits the classification framework, which enables the use of classification-based metrics that are more robust than the current CR metrics. Finally, MEI is also of practical use as it allows a user to search for all mentions of a particular entity or a group of entities of interest.
@@ 1 Introduction
The lack of domain generalization in CR models can largely be attributed to differences in annotation guidelines of popular CR benchmarks, specifically annotation guidelines about what constitutes a mention Porada et al. (2023). For example, OntoNotes Pradhan et al. (2013) does not annotate singletons, confounding mention identity with being referential. Thus, models trained on OntoNotes generalize poorly. The importance of mention detection for CR generalization is further highlighted by Gandhi et al. (2023), who show that solely annotating mentions is sufficient and more efficient for adapting pre-trained coreference models to new domains (in comparison to annotating coreference chains). Similarly, GPT-4 struggles with zero-/few-shot mention prediction, but given ground-truth mentions, its CR performance is competitive with the best-supervised models Le and Ritter (2023).
Given these observations, we hypothesize that current CR models, including large language models, generalize well at mention clustering but struggle to generalize on mention detection due to idiosyncrasies of different domains/benchmarks. We put forth an alternative formulation of the CR task where the entities of interest are provided as additional input. Assuming entities to be part of the input offloads the required domain adaptation from training to inference. Specifically, we propose the task of Major Entity Identification (MEI), where we assume the major entities of the narrative, defined as the most frequently occurring entities, to be provided as input along with the text (see Fig. 1). We focus on major entities for the following reasons: (a) Specifying major entities of a narrative is intuitively easier. (b) A handful of major entities often dominate any discourse. Table 1 shows that in LitBank roughly 6% of entities (490 of 7927) contribute to 60% of the mentions (16985 of 29103).
To test the generalizability of MEI, we adapt two literary CR benchmarks, namely LitBank Bamman et al. (2020) and FantasyCoref Han et al. (2021), and a state-of-the-art coreference model Toshniwal et al. (2021) to MEI. While there is a big gap in CR performance between in- and out-of-domain models Toshniwal et al. (2021), we show that this performance gap is much smaller for MEI (Section 5.1). To test this hypothesis further, we evaluate large language models (LLMs) for MEI in a few-shot learning setup. On CR, LLMs are shown to struggle with mention detection and perform worse than supervised models Le and Ritter (2023). Contrary to this, on MEI, top LLMs (e.g. GPT-4) are only slightly behind supervised models (Section 5.2). These experiments in the supervised setting and the few-shot setting demonstrate that the MEI task is more generalizable than CR.
@@
Additionally, we argue that MEI is easier to evaluate than CR. The MEI task can be viewed as a classification task in which any text span either refers to one of the input entities or the null class (minor entities and other non-mention spans). The classification formulation of MEI allows for the use of classification-based metrics that are more robust than the current CR metrics. Furthermore, MEI, by its definition, disregards insignificant and smaller clusters known to inflate the CR metrics Moosavi and Strube (2016); Lu and Ng (2020); Kummerfeld and Klein (2013). As an aside, formulating MEI as a classification task allows for a trivial parallelization across candidate spans (Appendix A.1).
@@ 3 Supervised MEI models
is performed using a Longformer-Large Beltagy et al. (2020), $\phi$, that we finetune for the task. Mentions (or spans) are encoded as $\mathbf{m}_{i}=\phi(m_{i},d)$ by concatenating the first, last, and an attention-weighted average of the token representations within the mention span. In MEI, an additional input is the set of designative phrases $\mathcal{P}$ for the major entities. Since each phrase is a derived from the document itself, we also obtain its encoding using the backbone: $\mathbf{e}_{j}=\phi(p(e_{j}),d)$.
@@ 4 Few-shot MEI with LLMs
CR or MEI can be addressed using separate few-shot prompting strategies for mention detection and mention clustering/identification. However, Le and Ritter (2023) found that this strategy faced significant challenges with mention detection, performing worse than a deterministic mention detector. Thus, they assume access to an oracle mention detector and focus their study to evaluating the linking capabilities of LLMs.
@@
We pose MEI as a structured generation task, prompting LLMs to reproduce documents and generate MEI tags at specific locations. Proprietary models like GPT-4 generally reproduce documents faithfully but for rare failures, we use the Needleman-Wunsch algorithm Needleman and Wunsch (1970) to align documents and extract tags In the case of open-source models, we employ regular expression-based constrained decoding with the outlines library Willard and Louf (2023)  [^2] https://outlines-dev.github.io/outlines/.
@@ 5 Experiments
Supervised models: Model hyperparameters are derived from Toshniwal et al. (2021). To ensure consistent performance across different numbers of target entities, we randomly select a subset of major entities at each training iteration (for more details, see Section A.2). All supervised models were trained five times with different random seeds, and we present aggregated results as the mean and standard deviation.
@@ 5.1 Experiments: Supervised Models
We train the fast-coref model Toshniwal et al. (2021) for CR and perform the following three inference-time adaptations for MEI:
Coref-ID: fast-coref uses active lists of entity representations, resolving coreference by associating mentions with existing clusters or generating new ones. During inference, we disable the cluster creation step and pre-fill the entity list with the encoded vector representations of the major entities. Hence, all the detected mentions either get mapped to one of the major entities or are discarded.
Coref-Cosine Map (Coref-CM): Since coreference clusters obtained from fast-coref lack explicit entity association, we employ the Kuhn-Munkres (KM) algorithm Munkres (1957) to find the optimal matching cluster for each major entity. The cost matrix uses the cosine similarity between the encoded representation of the major entities and that of the predicted cluster embeddings, both derived from fast-coref.
@@
MEIRa-S supports parallel batched processing since it does not update the working memory after associating mentions, i.e. the mentions need not be processed sequentially from left to right. Hence, post-mention detection (common to all models), MEIRa-S is about 25$\times$ faster than fast-coref when assessed across LitBank, FantasyCoref and AFT datasets on an NVIDIA RTX 4090 (see Fig. 3 in the appendix). Additionally, with the model’s small memory footprint during inference, the entire process can also be parallelized across chunks of documents making it extremely efficient. Hence, we pose MEIRa-S as a faster while competitive alternative to MEIRa-H (that requires dynamic updates and has similar computational performance as fast-coref).
@@
GPT-3.5 65.6 70.4 74.3 75.8
Code Llama-34B 63.4 70.8 68.3 72.7
Llama3-8B 50.5 57.8 46.3 52.1
Mistral-7B 62.1 71.1 61.2 70.9
@@ 5.2 Experiments: Few-shot prompting
We experiment with GPT-$4$Specifically, gpt-4-1106-preview OpenAI (2024), GPT-3.5555Specifically, gpt-3.5-turbo-1106, Code Llama-34B  Rozière et al. (2024), Mistral-7B  Jiang et al. (2023), and Llama3-8B. [^6] https://ai.meta.com/blog/meta-llama-3/ Following Le and Ritter (2023), we use the instruction-tuned versions for open-source models. These models were chosen for their ability to handle the extended context required for our benchmarks.
@@ 5.2.1 Linking Performance w/ Gold Mentions
The results in Table 5 show that GPT-4 surpasses the supervised MEIRa-H model in this setup. Among LLMs, GPT-4 is easily the best-performing model. Code Llama-34B performs the best among open-source models, closely followed by Mistral-7B. While Code Llama-34B is tailored for the code domain, surprisingly, it outperforms strong LLMs suited for natural language. This result corroborates a similar finding by  Le and Ritter (2023) for CR and related evidence regarding code pretraining aiding entity tracking  Kim et al. (2024). We find that Code Llama-34B performs close to GPT-3.5 for FantasyCoref, though a sizable gap remains for LitBank, potentially due to its linguistic complexity.
@@ 5.2.2 MEI Task Performance with LLMs
Code Llama-34B 19.4 23.4 19.4 16.2
Llama3-8B 29.2 32.8 24.5 27.1
Mistral-7B 28.0 30.9 14.9 15.3
@@
To further underscore the importance of mention detection, we also compare against the baseline GPT-4 w/ Ext det, which utilizes an external pre-trained mention detector followed by prompt-based linking (prompt in Table 11 of Appendix). We train the mention detector on the PreCo dataset Chen et al. (2018), which achieves a 93.8% recall and 53.1% precision on the combined FantasyCoref and LitBank validation sets. We observe that GPT-4 w/ Ext det is almost at par with the fully supervised MEIRa-H, again highlighting the strong mention linking capabilities of GPT-4.
Next, we present the results of our proposed Two-stage prompt, motivated by the Single prompt method’s failure with nested mentions. The first prompt asks GPT-4 to perform word-level MEI, by limiting the task to syntactic heads only. The second prompt then performs the task of mapping the identified syntactic heads to full mention spans. The results strongly validate our proposed approach with a relative improvement of more than 7% over the Single prompt method across all metrics and datasets. We also explore replacing the second step, i.e., head-to-span (H2S) retrieval, with an external tool. Specifically, we invert spaCy’s span-to-head mapping to obtain a head-to-span retriever. [^8] For the test set gold mentions of the two datasets, there were only two cases where spans had the same head. We handled these two cases manually.
@@
GPT-4 significantly improves in this setup, outperforming even the supervised model on LitBank. Given the strong performance of GPT-4 + spaCy H2S, we evaluate the open-source LLMs in only this setting. We observe a wide gap between GPT-4 and the open-source models. Llama3-8B surpasses other open-source models on both datasets, whereas the larger Code Llama-34B underperforms on the end-to-end task. This contrasts with the findings of the idealized golden mention setting, which assesses purely the model’s linking capabilities. The discrepancy between these results highlights the importance of evaluating in the realistic end-to-end setup.
@@ 6 Related Work
Character Identification
deals with specific characters from transcripts of TV shows and trains a model tailored to these constrained inputs Chen and Choi (2016); Zahiri and Choi (2017); Jiang et al. (2019).  Baruah and Narayanan (2023) introduced a dataset annotated with referent mentions of specific characters of interest. We differ from these works by adopting a generalized task formulation independent of annotation strategies and entity selection.
@@ 7 Conclusion
CR models are limited in their generalization capabilities owing to annotation differences and general challenges of domain adaptation. We propose MEI as an alternative to CR, where the entities relevant to the input text are provided as input along with the text. Our experiments demonstrate that MEI is more suited for generalization than CR. Additionally, MEI can be viewed as a classification task that (a) enables the use of more robust classification-based metrics and (b) a trivially parallelizable model across document spans, which gives a 25x speedup over a comparable coreference model, making MEI more suitable for longer narratives. Unlike CR, the formulation of MEI allows few-shot prompted LLMs to effectively compete with trained models. Our novel two-stage prompting and robust baseline methods empower top-performing LLMs like GPT-4 to achieve this. Our analysis indicates that this task holds promise for effectively evaluating the long-context referential capabilities of LLMs in an end-to-end manner.
@@ A.1 Linking Speed Comparison
This section compares the computational performance of fast-coref with the proposed MEIRa-S architecture. The classification formulation and the lack of an update step in MEIRa-S makes it a more efficient alternative to MEIRa-H and CR models. Fig. 3 displays the speed-up obtained in the identification module when assessed across documents with varying numbers of mentions. MEIRa-S consistently clocks a 20x efficiency across all ranges.
Figure 3: Linking speed comparison between  MEIRa-S and fast-coref for the combined LitBank and FantasyCoref test set. There exists 6 documents with (0, 100] mentions, 19 with (100, 500] mentions, 5 with (500, 1250] mentions and 3 with (1250, 2500] mentions.
@@ A.3 Prompts
1. Katharine Hilbery (#katharine_hilbery)
2. Mr. Denham (#mr._denham)
3. Mrs. Hilbery (#mrs._hilbery)
4. Mr. Hilbery (#mr._hilbery)
5. Mr. Fortescue (#mr._fortescue)

### v2

@@ Abstract
The limited generalization of coreference resolution (CR) models has been a major bottleneck in the task’s broad application. Prior work has identified annotation differences, especially for mention detection, as one of the main reasons for the generalization gap and proposed using additional annotated target domain data. Rather than relying on this additional annotation, we propose an alternative referential task, Major Entity Identification (MEI), where we: (a) assume the target entities to be specified in the input, and (b) limit the task to only the frequent entities. Through extensive experiments, we demonstrate that MEI models generalize well across domains on multiple datasets with supervised models and LLM-based few-shot prompting. Additionally, MEI fits the classification framework, which enables the use of robust and intuitive classification-based metrics. Finally, MEI is also of practical use as it allows a user to search for all mentions of a particular entity or a group of entities of interest.  [^1] Code for the paper is available at https://github.com/KawshikManikantan/MEI
@@ 1 Introduction
The lack of domain generalization in CR models can largely be attributed to differences in annotation guidelines of popular CR benchmarks, specifically annotation guidelines about what constitutes a mention Porada et al. (2023). For example, OntoNotes Pradhan et al. (2013) does not annotate singletons, confounding mention identity with being referential. Thus, models trained on OntoNotes generalize poorly Toshniwal et al. (2021). The importance of mention detection for CR generalization is further highlighted by Gandhi et al. (2023), showing that solely annotating mentions is sufficient and more efficient for adapting pre-trained CR models to new domains (in comparison to annotating coreference chains). Similarly, GPT-4 struggles with zero-/few-shot mention prediction, but with ground-truth mentions, its CR performance Le and Ritter (2023) is competitive with that of supervised models Toshniwal et al. (2021).
Given these observations, we hypothesize that current CR models, including large language models, generalize well at mention clustering but struggle to generalize on mention detection due to idiosyncrasies of different domains/benchmarks. We put forth an alternative referential task where the entities of interest are known and provided as additional input. Assuming entities to be part of the input offloads the required domain adaptation from training to inference. Specifically, we propose the task of Major Entity Identification (MEI), where we assume the major entities of the narrative, to be provided as input along with the text (see Fig. 1). We focus on major entities for the following reasons: (a) Specifying major entities of a narrative is intuitively easier. (b) A handful of major entities often dominate any discourse. Table 1 shows that in FantasyCoref roughly 16% of entities (942 of 5829) contribute to 63% of the mentions (35938 of 56968).
In this work, we adapt two literary CR benchmarks, namely LitBank Bamman et al. (2020) and FantasyCoref Han et al. (2021) by identifying frequently occurring entities as major entities and customizing a state-of-the-art coreference model Toshniwal et al. (2021) to MEI. Our tests for generalizability reveal that while there is a big gap in CR performance between in- and out-of-domain models Toshniwal et al. (2021), this performance gap is much smaller for MEI (Section 5.1). To test this hypothesis further, we evaluate large language models (LLMs) for MEI in a few-shot learning setup. On CR, LLMs are shown to struggle with mention detection and perform worse than supervised models Le and Ritter (2023). Contrary to this, on MEI, top LLMs (e.g. GPT-4) are only slightly behind supervised models (Section 5.2). These experiments in the supervised setting and the few-shot setting demonstrate that the MEI task is more generalizable than CR.
@@
Additionally, we argue that MEI is easier to evaluate than CR. The MEI task can be viewed as a classification task in which any text span either refers to one of the input entities or the null class (minor entities and other non-mention spans). The classification metrics maintain consistent granularity, proportionally penalize perturbations, and exhibit high discriminatory power while intuitively meeting multiple desired specifications (Moosavi and Strube, 2016; Recasens and Hovy, 2011).
Furthermore, MEI, by its definition, disregards insignificant and smaller clusters known to inflate the CR metrics Moosavi and Strube (2016); Lu and Ng (2020); Kummerfeld and Klein (2013). As an aside, formulating MEI as a classification task allows for a trivial parallelization across candidate spans (Appendix A.1).
@@ 3 Supervised MEI models
is performed using a Longformer-Large Beltagy et al. (2020), $\phi$, that we finetune for the task. Mentions (or spans) are encoded as $\mathbf{m}_{i}=\phi(m_{i},d)$ by concatenating the first, last, and an attention-weighted average of the token representations within the mention span. In MEI, an additional input is the set of designative phrases $\mathcal{P}$ for the major entities. Since each phrase is derived from the document itself, we also obtain its encoding using the backbone: $\mathbf{e}_{j}=\phi(p(e_{j}),d)$.
@@ 4 Few-shot MEI with LLMs
CR or MEI can be addressed using separate few-shot prompting strategies for mention detection and mention clustering/identification. However, Le and Ritter (2023) found that this strategy faced significant challenges with mention detection, performing worse than a deterministic mention detector. Thus, they assume access to an oracle mention detector and focus on evaluating LLMs’ linking capabilities.
@@
We pose MEI as a structured generation task, prompting LLMs to reproduce documents and generate MEI tags at specific locations. Proprietary models like GPT-4 generally reproduce documents faithfully but for rare failures, we use the Needleman-Wunsch algorithm Needleman and Wunsch (1970) to align documents and extract tags In the case of open-source models, we employ regular expression-based constrained decoding with the outlines library Willard and Louf (2023)
@@ 5 Experiments
Supervised models: Model hyperparameters are derived from Toshniwal et al. (2021). To ensure consistent performance across different numbers of target entities, we randomly select a subset of major entities at each training iteration (more details in Section A.2). Supervised models were trained five times with random seeds, and we present aggregated results as the mean and standard deviation.
@@ 5.1 Experiments: Supervised Models
We train the longdoc model Toshniwal et al. (2021) for CR and perform the following three inference-time adaptations for MEI:
Coref-ID: longdoc uses active lists of entity representations, resolving coreference by associating mentions with existing clusters or generating new ones. During inference, we disable the cluster creation step and pre-fill the entity list with the encoded vector representations of the major entities. Hence, all the detected mentions either get mapped to one of the major entities or are discarded.
Coref-Cosine Map (Coref-CM): Since CR clusters obtained from longdoc lack explicit entity association, we employ the Kuhn-Munkres (KM) algorithm Munkres (1957) to find the optimal matching cluster for each major entity. The cost matrix uses the cosine similarity between the encoded representation of the major entities and the predicted cluster embeddings, both derived from longdoc.
@@
MEIRa-S supports parallel batched processing since it does not update the working memory after associating mentions, i.e. the mentions need not be processed sequentially from left to right. Hence, post-mention detection (common to all models), MEIRa-S is about 25$\times$ faster than longdoc when assessed across LitBank, FantasyCoref and AFT datasets on an NVIDIA RTX 4090 (see Fig. 3 in the appendix). Additionally, with the model’s small memory footprint during inference, the entire process can also be parallelized across chunks of documents making it extremely efficient. Hence, we pose MEIRa-S as a faster while competitive alternative to MEIRa-H (that requires dynamic updates and has similar computational performance as longdoc).
@@
GPT-3.5 69.2 74.2 74.3 75.8
Code Llama-34B 67.0 72.4 68.9 73.1
Llama3-8B 53.8 60.6 50.2 53.4
Mistral-7B 67.3 75.8 61.6 73.9
@@ 5.2 Experiments: Few-shot prompting
We experiment with GPT-$4$Specifically, gpt-4-1106-preview OpenAI (2024), GPT-3.5555Specifically, gpt-3.5-turbo-0125, Code Llama-34B  Rozière et al. (2024), Mistral-7B  Jiang et al. (2023), and Llama3-8B. [^6] https://ai.meta.com/blog/meta-llama-3/ Following Le and Ritter (2023), we use the instruction-tuned versions for open-source models. These models were chosen for their ability to handle the extended context required for our benchmarks.
@@ 5.2.1 Linking Performance w/ Gold Mentions
The results in Table 5 show that GPT-4 surpasses the supervised MEIRa-H model in this setup. Among LLMs, GPT-4 is easily the best-performing model. Code Llama-34B performs the best among open-source models, closely followed by Mistral-7B. While Code Llama-34B is tailored for the code domain, surprisingly, it outperforms strong LLMs suited for natural language. This result corroborates a similar finding by  Le and Ritter (2023) for CR and related evidence regarding code pretraining aiding entity tracking  Kim et al. (2024). We find that Code Llama-34B performs close to GPT-3.5 for FantasyCoref, though a sizable gap persists in the Macro-F1 metric for LitBank , potentially due to its linguistic complexity.
@@ 5.2.2 MEI Task Performance with LLMs
Code Llama-34B 30.0 31.4 22.7 23.2
Llama3-8B 29.2 32.1 20.5 26.0
Mistral-7B 19.4 21.9 12.9 14.0
@@
To further underscore the importance of mention detection, we compare against GPT-4 w/ Ext det, which utilizes an external pre-trained mention detector followed by prompt-based linking (prompt in Table 11 of Appendix). We train the mention detector on the PreCo dataset Chen et al. (2018), which achieves a 93.8% recall and 53.1% precision on the combined FantasyCoref and LitBank validation sets. GPT-4 w/ Ext det performs at par with the fully supervised MEIRa-H, again highlighting the strong mention linking capabilities of GPT-4.
Next, we present the results of our proposed Two-stage prompt, motivated by the Single prompt method’s failure with nested mentions. The first prompt asks GPT-4 to perform word-level MEI, by limiting the task to syntactic heads only. The second prompt then performs the task of mapping the identified syntactic heads to full mention spans. The results strongly validate our proposed approach with a relative improvement of more than 10% over the Single prompt method across all metrics and datasets. We also explore replacing the second step, i.e., head-to-span (H2S) retrieval, with an external tool. Specifically, we invert spaCy’s span-to-head mapping to obtain a head-to-span retriever. [^8] For the test set gold mentions of the two datasets, there were only two cases where spans had the same head. We handled these two cases manually.
@@
GPT-4 significantly improves in this setup, outperforming even the supervised model on LitBank. Given the strong performance of GPT-4 + spaCy H2S, we evaluate the open-source LLMs in only this setting. We observe a wide gap between GPT-4 and the open-source models. Llama3-8B outperforms other open-source models on both datasets in Micro-F1 and stays competitive with the larger Code Llama-34B in Macro-F1. However, this contrasts with Llama3-8B’s significant lag in the idealized golden mention setting, which solely evaluates the model’s linking capabilities.
@@ 6 Related Work
Entity-centric tasks
similar to MEI include character identification, where either annotations are restricted to a subset of entities Baruah and Narayanan (2023) or custom models are developed to extract mentions of specific characters from TV show transcripts Chen and Choi (2016); Zahiri and Choi (2017). We differ from these works by adopting a generalized task formulation independent of annotation strategies and entity selection. Another task, Entity Linking Ji et al. (2015) extracts distinct entities from a document and links them to external Knowledge Bases. In contast, MEI focuses on retrieving all mentions (including nominals and pronominals) of a specific set of key entities, extracted solely from the context of the document.
@@ 7 Conclusion
CR models are limited in their generalization capabilities owing to annotation differences and general challenges of domain adaptation. We propose MEI as an alternative to CR, where the entities relevant to the input text are provided as input along with the text. Our experiments demonstrate that MEI is more suited for generalization than CR. Additionally, MEI can be viewed as a classification task that enables the use of intuitive metrics. A trivially parallelized variation (MEIRa-S), gives a 25x speedup over a comparable CR model, making it more suitable for longer narratives. Unlike CR, the formulation of MEI allows few-shot prompted LLMs to effectively compete with trained models. Our novel two-stage prompting and robust baseline methods empower top-performing LLMs like GPT-4 to achieve this. Our analysis indicates that this task holds promise for effectively evaluating the long-context referential capabilities of LLMs in an end-to-end manner.
Potential applications of MEI include domains such as film and literature, where metadata about salient entities can be sourced from external databases like IMDb or SparkNotes. Additionally, MEI can be applied to the analysis of documents like of financial and legal reports, when the user is familiar with the relevant entities. Lastly, recent research Lin and Zeldes (2024) indicates that LLMs can assist or automate the extraction of salient entities, a direction we intend to explore in future work.
@@ A.1 Linking Speed Comparison
This section compares the computational performance of longdoc with the proposed MEIRa-S architecture. The classification formulation and the lack of an update step in MEIRa-S makes it a more efficient alternative to MEIRa-H and CR models. Fig. 3 displays the speed-up obtained in the identification module when assessed across documents with varying numbers of mentions. MEIRa-S consistently clocks a 20x efficiency across all ranges.
Figure 3: Linking speed comparison between  MEIRa-S and longdoc for the combined LitBank and FantasyCoref test set. There exists 6 documents with (0, 100] mentions, 19 with (100, 500] mentions, 5 with (500, 1250] mentions and 3 with (1250, 2500] mentions.
@@ A.3 Prompts
1. Katharine Hilbery - #1
2. Mr. Denham - #2
3. Mrs. Hilbery - #3
4. Mr. Hilbery - #4
5. Mr. Fortescue - #5